OpenStudy (anonymous):

HELP!!! Give an example of a logarithm that you would need to use the change of base formula to evaluate. Why do you need to use the change of base formula?

OpenStudy (campbell_st):

what about logs with different bases.... use change of base to change the left hand side to base 5.. \[\log _{3} 2x = \log _{5} x + 7\]

OpenStudy (anonymous):

confused still...

OpenStudy (campbell_st):

ok... so you can't solve for x until you get the log bases the same.... you need change of base on the left hand side \[\frac{\log _{5}(x)}{\log _{5}(2)} = \log _{5}(x + 7)\]

OpenStudy (anonymous):

is there a more simple example? like could you say..hold on

OpenStudy (anonymous):

\[\log_{5} 3\]

OpenStudy (anonymous):

=

OpenStudy (anonymous):

log3/log5

OpenStudy (anonymous):

@campbell_st ?

OpenStudy (campbell_st):

what about \[\log _{3} 2x = \log _{7}(4)\]

OpenStudy (anonymous):

does it have to have a variable?